OpenStudy (anonymous):

given that x=3 is a zero of P(x) =x^3-5x^2-2x+24 find all other zeros of p(x)

OpenStudy (anonymous):

idea is to factor out \[x^3-5x^2-2x+24=(x-3)\times q(x)\]

OpenStudy (anonymous):

q(x) will be a polynomial of degree 2 and you find the zeros either by factoring or using quadratic formula. you find q(x) by dividing. synthetic division is easiest

OpenStudy (anonymous):

but really easiest is using a calculator. http://www.wolframalpha.com/input/?i=P%28x%29+%3Dx^3-5x^2-2x%2B24

OpenStudy (amistre64):

synthetic division :) coeffs of ploy addends ----------- x | new polys

OpenStudy (amistre64):

1 -5 -2 24 0 3 -6 -24 ------------ 3 | 1 -2 -8 0 ; the zero at the end tells us its a zero of the poly x^2 -x -8 is whats left to get zeros for, and quadratics thed to be easier to deal with;

OpenStudy (amistre64):

opps... i dropped out the 2, this is better x^2 -2x -8

OpenStudy (amistre64):

we can either factor this, or use the quadratic formula to determine the zeros; the quad form will always provide results whereas factoring depends on luck

OpenStudy (amistre64):

(1/2) (2+-sqrt(4-4(1)(-8))) (1/2) (2+-sqrt(4+32)) (1/2) (2+-sqrt(36)) (1/2) (2+- 6) (1/2) (-4 or 8) means x = -2 or 4